Calculate the molar concentration of solute in the following:

  1. 7.59g of NH3 in 1.500L solution.
  2. 0.291g of H2SO4 in 8.65L solution.
  3. 85.2g of NaOH in 5.00L solution.
  4. 31.25g of NaCl in 2.00L solution.
  5. 6mg of atmospheric N2 in water with volume of 2.00L.
1
Expert's answer
2021-11-02T02:07:11-0400

Molarity of "NH_3=[\frac{7.59}{17.031}]÷1.500=0.297M"


"H_2SO_4=[\frac{0.291}{
98.079 }]÷8.65=0.00034M"


"NaOH=[\frac{85.2}
{
39.997 
}]÷5.00=0.43M"


"NaCl=[\frac{31.25}{
58.44 
}]÷2.00=0.267M"